Allison:

Given the function g(x)=-x^2+7x+14g(x)=−x 2 +7x+14, determine the average rate of change of the function over the interval 1\le x \le 101≤x≤10.

dude:

Between 1 and 10?

Allison:

Yeah not sure why it copies as 101

Allison:

-x^2+7x+14

dude:

Formatting dw \(-(1)^2+7(1)+14\) <-- this should be ez \(=-(10)^2+7(10)+14\)

Allison:

19 and 64

Allison:

So 19 and 61... 1 and 10

dude:

No no 20 and -16 But same idea \(\dfrac{-16-20}{10-1}\)

Allison:

Where you get those from? My calculator says 19 and 64

dude:

\(=-(1)^2+7(1)+14\) \(-1+7+14=20\) ~~~~~~~~~ \(=-(10)^2+7(10)+14\) \(-100+70+14\) \(-30+14=-16\)

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude No no 20 and -16 But same idea \(\dfrac{-16-20}{10-1}\) \(\color{#0cbb34}{\text{End of Quote}}\)